# for Scalable MoE Pruning

### v1

@@ 4.3 Towards $O(n)$: Probabilistic Interpretation
$\sum_{i\in C_{l}}\sum_{l}\|W_{i}-\mu_{l}\|^{2}$ (9)
Practically, we found that the agglomerative clustering algorithm Sneath and Sokal (1973) performs well. [^3] We tried other clustering algorithms in the Appendix. Specifically, clusters are initialized as individual experts and then iteratively merged, with a termination condition that prevents the experts within each cluster from being too dissimilar. This condition is tuned based on the desired pruning ratio.
Lastly, if we allow inference on some data, we can improve Eq. 8 with coactivation statistics $a_{i,j}$, which measure the frequency with which $E_{i},E_{j}$ are selected simultaneously. [^4] We normalize $a_{i,j}$ by dividing it with the total coactivations in one layer. However, these coactivation statistics depend on the given data, whose distribution may differ from the test data. Therefore we combine the two as follows:
$b_{i,j}=-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$ (10)
We recap the algorithm in the Appendix (Alg 1).
4.4 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
1st-order Taylor Approximation
While the $O(n)$ approach immensely reduces the cost to obtain the probability distribution, we can further reduce the number of forward passes needed by precalculating $P(E_{i})$s in Eq. 7.
$\sum_{i}\mathcal{E}_{i}<\sum_{i}\gamma\|\theta_{i}-\theta_{C}\|^{2}$ (12)
Selective Reconstruction of Experts
To balance these two types of errors, we selectively decide whether to reconstruct. We observe that $\sum_{i}\mathcal{E}_{i}$ increases if the total number of clusters in a layer decreases, as this would introduce more $\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ terms. Therefore, if the total number of clusters is below a threshold $\kappa$, we use $\theta_{C}=\overline{\theta_{i}}$ to minimize $\sum_{i}\mathcal{E}_{i}$. Otherwise, we set $\theta_{C}$ as the expert within the cluster $\{\theta_{i}\}_{i=1}^{|C|}$ closest to the $\overline{\theta_{i}}$, to minimize $\mathcal{E}_{d}$. The router weight reconstruction is done similarly, following its corresponding expert. The final algorithm is summarized in the Appendix (Alg 2).
5 Robustness of Structured-Then-Unstructured Pruning
In this section, we deliver the key insights into why STUN works: MoEs are inherently robust to expert pruning, and the expert-pruned MoE is robust to subsequent post-hoc unstructured pruning.
MoE Robustness to Structured (Expert) Pruning
To support this, we find a resemblance between the MoE (Eq. 3) and the targeted dropout Gomez et al. (2019). Targeted dropout achieves a higher post-hoc pruning ratio by stochastically dropping out unimportant units or weights during training. Formally, in each training step, $M(x;\theta)$ is modified as follows:
$M(x;\theta-U(\theta)\odot D)$ (13)
where $U(\theta)$ is a function that extracts only an unimportant part of a given weight $\theta$, and $D$ is a stochastic mask to ensure $U(\theta)$ also has a chance to be updated. Gomez et al. (2019) designed $U(\theta)$ as the bottom $|\theta|-k$ elements sorted by their absolute values, and $D\sim Bernoulli(\alpha)$ where $\alpha$ is the dropout ratio. After training, the performance is better retained when $U(\theta)$ is pruned.
We recognize the resemblance between the MoE (Eq. 3) and the targeted dropout (Eq. 13). If we interpret $U(\theta)$ as $\bigcup_{i\notin\mathcal{T}}E(\theta_{i})$, then $M(x;\theta-U(\theta))$ resembles Eq. 3. Moreover, just as $D$ is added in Eq. 13 to ensure that $U(\theta)$ is updated, MoE ensures that $U(\theta)$ is updated through the stochastic nature of $r(x)$.
Therefore, after training, MoE is expected to be robust to pruning $U(\theta)$, which, in this context, corresponds to pruning some experts $E(\theta_{i})$s. However, identifying $U(\theta)$ is not as straightforward as in Gomez et al. (2019). In MoE training, $U(\theta)$ is conditioned on each token, such that we need to marginalize token dependency, by identifying unimportant experts across a broad range of inputs. This challenge is connected to the problem of finding $S$ in Eq. 4, which minimizes the reconstruction loss.
Expert-Pruned MoE Robustness to Unstructured Pruning
Robustness to unstructured pruning can be estimated by the kurtosis of weights Mason-Williams and Dahlqvist (2024). Kurtosis is expressed as follows:
In contrast, coarse structured pruning, such as expert pruning, is less likely to decrease the kurtosis value, since the assumption $\theta\sim\mathcal{N}$ still holds for remaining experts. This implies that expert pruning preserves the robustness of unstructured pruning, unlike applying unstructured pruning with a similar pruning ratio.
6 Experiments
model sparsity method GSM8K Avg(→) ARC-c ARC-e HellaSwag MMLU
Arctic 0% unpruned 70.74 68.33 56.91 84.60 66.94 64.86
40% STUN (w/ OWL) 70.28 67.66 57.68 83.29 64.94 64.75
OWL 63.76 67.35 56.74 84.13 65.08 63.47
STUN (w/ Wanda) 69.60 67.64 57.25 83.63 64.86 64.81
Wanda 64.59 67.54 57.00 84.64 65.19 63.32
65% STUN (w/ OWL) 43.97 62.67 51.54 80.01 59.91 59.24
OWL 13.42 56.68 44.37 76.64 53.69 52.02
OWL 1.29 45.20 24.15 49.79 49.27 57.60
OWL 19.64 57.74 45.48 76.60 52.47 56.42
25% Ours $O(1)$ 59.30 85.44 88.13 64.42 64.52 35.40 71.84 75.37 68.05
Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 58.19 84.89 87.34 65.24 62.47 35.60 70.04 75.85 67.45
25% Ours $O(1)$ 52.73 81.82 83.09 60.84 63.34 31.60 68.59 72.69 64.34
Lu et al. (2024) $O(\frac{k^{n}}{\sqrt{n}})$ 51.62 81.94 83.64 61.60 58.72 33.00 67.87 75.37 64.22
(a) Arctic (128x3.66B)
(b) Mixtral-8x7B
(c) Mixtral-8x22B
Figure 2: GSM8K 5-shot accuracy comparing STUN and unstructured pruning for various MoEs.
In this section, we aim to address the following research questions:
  • RQ1: Does STUN outperform unstructured pruning alone?
  • RQ2: Does the expert pruning we propose outperform existing baselines?
  • RQ3: Does STUN favor recent MoE trends with a large number of small experts?
  • RQ4: Is each component of STUN essential for achieving performance gains?
  • RQ5: Does STUN generalize to non-MoE models as well?
6.1 Experimental Settings
We use Snowflake Arctic Snowflake (2024) as a representative large MoE, with a total of 480B parameters and 128 experts. To compare our method with previous works Lu et al. (2024), we also experiment with Mixtral Jiang et al. (2024).
Tasks and Datasets
In contrast to previous unstructured pruning studies Sun et al. (2024); Yin et al. (2024), we also evaluate on the NLG task, GSM8K Cobbe et al. (2021), where maintaining performance proves to be much more challenging (see Table 1). We further assess performance on four NLU tasks– ARC-challenge and ARC-easy Clark et al. (2018), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2021). When comparing with expert pruning methods, following previous work Lu et al. (2024), we also conduct a zero-shot evaluation on BoolQ Wang et al. (2019), OpenBookQA Mihaylov et al. (2018), RTE Wang et al. (2018), WinoGrande Sakaguchi et al. (2021).
To provide some data for inference, we employ the C4 dataset Raffel et al. (2020), following the baselines Yin et al. (2024); Sun et al. (2024); Lu et al. (2024).
Implementation Details
We explore the values of $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for Snowflake Arctic, the largest MoE in our experiments, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, meaning no GPU calls are required for expert pruning. We evaluate methods using the lm-evaluation-harness Gao et al. (2021). Due to the model size, we use 4-bit quantization Dettmers et al. (2023) for experiments with Mixtral-8x22B and Arctic. The expert pruning ratios are set to 20%, 12.5%, and 10% for Snowflake Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively. We use $\kappa=3$ for selective reconstruction. More detailed implementation information is provided in the Appendix.
6.2 Experimental Results
6.2.1 RQ1: STUN Outperforms Unstructured Pruning
For example, with 40% of sparsity for the Arctic, STUN neatly retains the original GSM8K performance, while unstructured pruning results in a noticeable performance drop. This is consistent for different unstructured pruning methods, Wanda, as well. For 65% of sparsity for Mixtral-8x7B-Instruct, STUN’s GSM8K performance is nearly 20 times better than that of unstructured pruning. In the ARC-challenge, the unstructured pruning performance falls below the random-guess accuracy of 25.02 Clark et al. (2018), whereas STUN maintains a significantly higher performance, achieving twice the score.
6.2.2 RQ2: Our $O(1)$ Expert Pruning Outperforms Existing $O(\frac{k^{n}}{\sqrt{n}})$
Table 2 shows that our proposed $O(1)$ expert pruning method is highly effective, outperforming the previous $O(\frac{k^{n}}{\sqrt{n}})$ solution. This is because we derive the latent structure from the pretrained MoE, while the previous work Lu et al. (2024) relies solely on the given calibration data. This validates our design of $O(1)$ in section 4.
6.2.3 RQ3: STUN Favors Large Number of Small Experts
Figure 2 illustrates the trend of STUN in 3 different MoEs. The performance gap between STUN and unstructured pruning increases as the MoE has more experts with small sizes (from (c) to (a)). This is because having more experts, rather than having fewer but larger ones, provides greater flexibility to our expert pruning. Notably, MoEs with a large number of small experts are favored in recent models He (2024).
Cluster (§$\S$ 4.3) Reconstruct (§$\S$ 4.4) LM-eval Avg
Ours Ours 59.58
DSatur Ours 58.59
Ours Always 57.60
Ours Never 59.22
6.2.4 RQ4: Ablation
To validate our design of expert pruning in sections 4.3 and 4.4, we evaluate alternative approaches to expert-prune Mixtral-8x7B at 50% sparsity. Table 3 confirms that our design choices are valid. Our agglomerative clustering algorithm outperforms the DSatur algorithm, an alternative clustering algorithm we discuss in the Appendix. Additionally, selective reconstruction proves superior to always or never reconstructing, as shown in the last two rows. Detailed evaluation results are provided in the Appendix.
6.2.5 RQ5: STUN Outperforms Unstructured Pruning in non-MoEs
To investigate whether STUN is generalizable to non-MoE as well, we employ a state-of-the-art structured pruning algorithm for non-MoE model, namely, LLM-surgeon van der Ouderaa et al. (2024) with 5% sparsity before performing unstructured pruning, which is OWL in our case. Figure 3 illustrates that such STUN outperforms unstructured pruning.
(a) Llama-2 7B
(b) Llama-2 13B
7 Conclusion
In this paper, we proposed STUN for MoEs, which outperforms unstructured-only pruning. Although this may seem counterintuitive, we provide both theoretical and empirical evidence demonstrating why designing expert pruning before unstructured pruning is beneficial. Since existing expert pruning requires $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, we proposed $O(1)$ complexity, which outperforms the more expensive approaches. Our experimental results show STUN  is highly effective– For 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to.
Limitation
Since our method utilizes unstructured pruning in the second stage, we share the same disadvantages with unstructured pruning, that is, it may need specialized hardware for acceleration. However, it is shown that general-purpose hardware, such as CPU, can successfully accelerate unstructure-pruned networks NeuralMagic (2021). Therefore we have the same potential as existing unstructured pruning methods– with the support from the community, our method will scale up to existing accelerators.
Appendix
Algorithm of $O(1)$ Expert Pruning
Algorithms 1 and 2 describe the expert pruning we proposed. Note that we do not introduce any GPU inference in both two algorithms.
@@ 7 Conclusion
$t\leftarrow$ Threshold to determine pruning ratio

### v2

@@ 4.2.2 Towards $O(n)$: Probabilistic Interpretation
which can be improved with coactivation statistics $a_{i,j}$, if we allow some inference cost. As a result, we illustrate our clustering algorithm in Alg 1, whose detailed derivation can be found in Appendix A.1.
@@
$t\leftarrow$ Threshold to determine sparsity
$c\leftarrow$ The mapping from expert to cluster of the similar experts
@@
4.2.3 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
While the previous section immensely reduces the cost to obtain the probability distribution to $O(n)$ by requiring only $P(E_{i})$s, we can further reduce the number of forward passes– We aim to remove the GPU calls for $P(E_{i})$, which is needed as in Eq. 10.
The key idea is approximating $E_{i}$’s reconstruction loss value $\mathcal{E}_{i}=\|M(x;\theta)-M(x;\theta-\theta_{i})\|_{F}$. To address this, with 1st order Taylor approximation, we find the expert closest to $\bar{\theta_{i}}$ within each cluster has the highest priority to be retained. We assign ranks similarly, and the same greedy algorithm is applied to optimize Eq. 5. Additionally, we selectively reconstruct the expert. The final algorithm is summarized in Alg 2, which is described in detail in Appendix A.2.
@@ 4.2.3 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
4.3 Unstructured Pruning on Expert-pruned Model
Our main conjecture for STUN is intra-expert sparsity yet remains intact after the first phase. Specifically, we propose to pursue fine-grained sparsity within the remaining experts, by leveraging unstructured pruning methods designed for general LLM , such as OWL Yin et al. (2024) or Wanda Sun et al. (2024).
Now we theoretically verify our conjecture: intra-expert sparsity remains high, or, even higher, after the expert pruning phase. In other words, we explain why performing unstructured pruning after expert pruning is better than continuously performing unstructured pruning only.
Intra-expert sparsity, formally speaking, robustness to unstructured pruning, can be estimated by the kurtosis of weights Mason-Williams and Dahlqvist (2024). Kurtosis is expressed as follows:
In contrast, coarse structured pruning, such as expert pruning, is less likely to decrease the kurtosis value, since the assumption $\theta\sim\mathcal{N}$ still holds for remaining experts. This implies that expert pruning preserves the robustness of unstructured pruning, unlike applying unstructured pruning with a similar sparsity. [^5] In our experiments, the kurtosis increased from 14248 to 15623 after expert pruning.
5 Experiments
model sparsity method GSM8K Avg(→) ARC-c ARC-e HellaSwag MMLU
Arctic 0% unpruned 70.74 68.33 56.91 84.60 66.94 64.86
40% STUN (w/ OWL) 70.28 67.66 57.68 83.29 64.94 64.75
OWL 63.76 67.35 56.74 84.13 65.08 63.47
STUN (w/ Wanda) 69.60 67.64 57.25 83.63 64.86 64.81
Wanda 64.59 67.54 57.00 84.64 65.19 63.32
65% STUN (w/ OWL) 43.97 62.67 51.54 80.01 59.91 59.24
OWL 13.42 56.68 44.37 76.64 53.69 52.02
OWL 1.29 45.20 24.15 49.79 49.27 57.60
LLM-Pruner 1.29 31.74 22.78 45.96 35.09 23.13
OWL 19.64 57.74 45.48 76.60 52.47 56.42
model sparsity method cost Avg
Mixtral-8x7B (Instruct) 0% unpruned 69.98
25% Ours $O(1)$ 68.05
Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 67.45
Mixtral-8x7B 0% unpruned 67.58
25% Ours $O(1)$ 64.34
Lu et al. (2024a) $O(\frac{k^{n}}{\sqrt{n}})$ 64.22
Table 3: Comparing the average performance of 8 tasks of the proposed expert pruning, with other baselines.
sparsity ARC-C BoolQ HellaSwag MMLU RTE WinoGrande Avg
Unpruned 0% 59.4 84.2 84.0 67.9 70.4 75.6 71.5
Ours 25% 55.6 83.1 81.1 63.3 68.6 72.7 70.7
SEER-MoE Muzio et al. (2024) 25% - - - 56.7 - - -
Expert Drop He et al. (2024) 25% 53.2 77.7 80.5 52.2 55.6 76.8 66.0
Layer Drop He et al. (2024) 25% 47.7 85.3 75.2 67.3 69.7 74.6 70.0
Table 4: Comparison of efficient expert pruning methods on Mixtral-8x7B.
sparsity cost GSM8K
unpruned 0% 63.46
Ours 25% $O(1)$ 53.22
Lu et al. (2024a) 25% $O(\frac{k^{n}}{\sqrt{n}})$ 48.52
Table 5: GSM8K accuracy comparison with baseline on Mixtral-8x7B-Instruct.
5.1 Experimental Settings
We use Snowflake Arctic Snowflake (2024) as a representative large MoE, with a total of 480B parameters and 128 experts. To compare our method with previous works Lu et al. (2024a), we also experiment with Mixtral Jiang et al. (2024).
Tasks and Datasets
In contrast to previous unstructured pruning studies Sun et al. (2024); Yin et al. (2024), we also evaluate the NLG task, GSM8K Cobbe et al. (2021), where maintaining performance proves to be much more challenging (see Table 2; Appendix F). We further assess performance on four NLU tasks– ARC-challenge and ARC-easy Clark et al. (2018), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2021). When comparing with expert pruning methods, following previous work Lu et al. (2024a), we also conduct a zero-shot evaluation on BoolQ Wang et al. (2019), OpenBookQA Mihaylov et al. (2018), RTE Wang et al. (2018), WinoGrande Sakaguchi et al. (2021).
To provide some data for inference, we employ the C4 dataset Raffel et al. (2020), following the baselines Yin et al. (2024); Sun et al. (2024); Lu et al. (2024a).
@@ 5.1 Experimental Settings
We explore the values of $(\lambda_{1},\lambda_{2})\in\{(0,1),(1,0),(1,1)\}$, except for Snowflake Arctic, the largest MoE in our experiments, where we only consider $(\lambda_{1},\lambda_{2})=(1,0)$, meaning no GPU calls are required for expert pruning. The sparsity for expert-level structured pruning is set to 20%, 12.5%, and 10% for Snowflake Arctic, Mixtral-8x7B, and Mixtral-8x22B respectively. We use $\kappa=3$ for selective reconstruction. More detailed implementation information and hyperparameter decisions are provided in Appendix.
5.2 Experimental Results
5.2.1 RQ1: STUN Outperforms Unstructured Pruning
For example, with 40% of sparsity for the Arctic, STUN neatly retains the original GSM8K performance, while unstructured pruning results in a noticeable performance drop. This is consistent for different unstructured pruning methods, Wanda, as well. As the sparsity increases, the table shows that STUN can maintain the original performance much better than the baselines– For 65% sparsity, STUN’s GSM8K performance is nearly 30%p better than that of unstructured pruning. With different models, we observe different gaps but a similar trend– For 65% of sparsity for Mixtral-8x7B-Instruct, STUN’s GSM8K performance is nearly 20 times better than that of unstructured pruning. In the ARC-challenge, the unstructured pruning performance falls below the random-guess accuracy of 25.02 Clark et al. (2018), whereas STUN maintains a significantly higher performance, achieving twice the score. Mixtral-8x22B shows a similar trend.
sparsity $O(n)$ $O(1)$
25% 63.97 64.34
50% 59.90 59.58
Table 6: Comparing the average performance of 8 tasks of our $O(n)$ and $O(1)$ algorithms for expert pruning on Mixtral-8x7B.
(a) Arctic (128x3.66B)
(b) Mixtral-8x7B
(c) Mixtral-8x22B
(d) OLMoE-7B
Figure 3: Comparing STUN and unstructured pruning for various MoEs.
(e) Llama-2 7B
(f) Llama-2 13B
5.2.2 RQ2: Our $O(1)$ Expert Pruning Outperforms Existing Methods
Tables 3 and 5 show that our proposed $O(1)$ expert pruning method is highly effective, outperforming the previous $O(\frac{k^{n}}{\sqrt{n}})$ solution. This is because we derive the latent structure from the pretrained MoE, while the previous work Lu et al. (2024a) relies solely on the given calibration data. This validates our design of $O(1)$ in section 4. The detailed results are in Appendix E.
Table 4 shows that our expert pruning outperforms other efficient pruning methods. Compared to SEER-MoE, our pruning clearly outperforms the performance in MMLU, by a substantial margin. Moreover, ours achieves a higher average performance compared to the results reported by He et al. (2024). [^6] Note that SEER-MoE Muzio et al. (2024) only reveals the performance of MMLU, and we used different metrics following He et al. (2024).
Moreover, Table 6 shows that our $O(1)$ expert pruning method achieves similar performance to our $O(n)$ method. This supports our choice to use the $O(1)$ method, which is more efficient.
Train GPU cost GPUs #
Lu et al. (2024a) 0 infeasible [^7] In detail, 23951146041928082866135587776380551750 forward passes per layer at minimum. > 8
STUN (w/ OWL) 0 1.12h 1
OWL Yin et al. (2024) 0 1.12h 1
STUN (w/ Wanda) 0 0.58h 1
Wanda Sun et al. (2024) 0 0.58h 1
Table 7: Cost comparison of diverse pruning methods with Snowflake-Arctic. Train: training cost, GPU #: number of GPUs required for pruning.
5.2.3 RQ3: STUN Adapts Large Number of Small Experts
Figure 3 illustrates the trend of STUN in different MoEs. The performance gap between STUN and unstructured pruning increases as the MoE has more experts with small sizes (from (c) to (a)). This is because having more experts, rather than having fewer but larger ones, provides greater flexibility to our expert pruning. Notably, MoEs with a large number of small experts are favored in recent models He (2024).
5.2.4 RQ4: STUN Outperforms Unstructured Pruning in non-MoEs
To investigate whether STUN is generalizable to non-MoE as well, we employ a state-of-the-art structured pruning algorithm for non-MoE models, namely, LLM-surgeon van der Ouderaa et al. (2024) with 5% sparsity before performing unstructured pruning, which is OWL in our case. Figure 4 illustrates that such STUN outperforms unstructured pruning.
6 Cost Analysis
Table 7 shows the cost comparison between diverse pruning methods. While none of the pruning methods require training cost, Lu et al. (2024a) is infeasible due to its prohibitive number of forward passes in GPUs. Due to the efficiency of proposed expert pruning, STUN is as efficient as the unstructured pruning method it uses, making it a feasible pruning method even for large MoEs, such as Snowflake-Arctic.
7 Conclusion
In this paper, we proposed STUN– an interpolation between structured and unstructured pruning, leveraging both inter- and intra-expert sparsity.We provide both theoretical and empirical evidence demonstrating why designing expert pruning before unstructured pruning is beneficial.
Limitation
Since our method utilizes unstructured pruning in the second stage, we share the same disadvantages with unstructured pruning, that is, on some hardware, the acceleration may not be trivial. However, it is shown that some hardware, such as CPU, can successfully accelerate unstructure-pruned networks NeuralMagic (2021), and ongoing research is actively developing methods to achieve similar speedups on GPUs Mishra et al. (2021); Zhao et al. (2024). With a substantial body of work dedicated to unstructured pruning Frantar and Alistarh (2023); Das et al. (2024); Dong et al. (2024); Sun et al. (2024); Yin et al. (2024); Li et al. (2024); Hu et al. (2024); Lu et al. (2024b), we believe this approach remains highly relevant and leads to practical performance improvements with ongoing hardware support.
We already have shown our method works in versatile extreme settings, such as non-MoE models, or high sparsity, etc. We leave more extensive evaluations, such as performance under highly imbalanced or skewed data distributions, as future work.
Appendix A Derivation From $O(\frac{k^{n}}{\sqrt{n}})$ to $O(1)$
A.1 Towards $O(n)$: Probabilistic Interpretation
As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ (Eq. 4) as:
$\textrm{argmax}_{S}P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ (9)
where $s_{i}$s are the experts included in the expert set $S$, and $P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ is the joint probability of pruning $S$. One intuitive way to design $P$ so that Eq. 9 yields the same $S$ to minimize Eq. 4 is as follows:
$P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})=\frac{1}{Z}\cdot\frac{1}{\mathcal{E}_{S}}$ (10)
where $Z$ is the normalization factor. Each joint probability needs $O(1)$ GPU calls, since $\mathcal{E}_{S}$ needs the output of $M(x;\theta-\theta_{S})$.
Section 4.2.1 corresponds to enumerating joint probability from all combinations, requiring ($n\choose{|S|}$ different values, which is compute-intensive. When chain rule is applied, Eq. 9 can be reformulated as follows:
Our contribution is greedy optimization without compromise for Eq. 11– We decompose the multiplication of Eq. 11 at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability. For simplicity, we will omit $X_{k}$s from this point on.
Clustering the Similar Experts
$b_{i,j}=-\|W_{i}-W_{j}\|_{F}$ (13)
$\sum_{i\in C_{l}}\sum_{l}\|W_{i}-\mu_{l}\|^{2}$ (14)
Practically, we found that the agglomerative clustering algorithm Sneath and Sokal (1973) performs well. [^8] We tried other clustering algorithms in the Appendix. Specifically, clusters are initialized as individual experts and then iteratively merged, with a termination condition that prevents the experts within each cluster from being too dissimilar. This condition is tuned based on the desired sparsity.
Lastly, if we allow inference on some data, we can improve Eq. 13 with coactivation statistics $a_{i,j}$, which measure the frequency with which $E_{i},E_{j}$ are selected simultaneously. [^9] We normalize $a_{i,j}$ by dividing it with the total coactivations in one layer. However, these coactivation statistics depend on the given data, whose distribution may differ from the test data. Therefore, we combine the two as follows:
$b_{i,j}=-\lambda_{1}\|W_{i}-W_{j}\|_{F}+\lambda_{2}a_{i,j}$ (15)
We recap the algorithm in the Appendix (Alg 1).
A.2 Towards $O(1)$: Taylor Approximation and Selective Reconstruction
1st-order Taylor Approximation
While previous section immensely reduces the cost to obtain the probability distribution to $O(n)$ by requiring only $P(E_{i})$s, we can further reduce the number of forward passes– We aim to remove the GPU calls for $P(E_{i})$, which is needed as in Eq. 10.
$\sum_{i}\mathcal{E}_{i}<\sum_{i}\gamma\|\theta_{i}-\theta_{C}\|^{2}$ (17)
Selective Reconstruction of Experts
To balance these two types of errors, we selectively decide whether to reconstruct. We observe that $\sum_{i}\mathcal{E}_{i}$ increases if the total number of clusters in a layer decreases, as this would introduce more $\|E^{\prime}(\theta_{i})\cdot(\theta_{i}-\theta_{C})\|^{2}$ terms. Therefore, if the total number of clusters is below a threshold $\kappa$, we use $\theta_{C}=\overline{\theta_{i}}$ to minimize $\sum_{i}\mathcal{E}_{i}$. Otherwise, we set $\theta_{C}$ as the expert within the cluster $\{\theta_{i}\}_{i=1}^{|C|}$ closest to the $\overline{\theta_{i}}$, to minimize $\mathcal{E}_{d}$. The router weight reconstruction is done similarly, following its corresponding expert.
Appendix B Implementation Details